# Computation of attractor dimension and maximal sums of Lyapunov exponents using polynomial optimization

### v1

@@ Abstract
Two approaches are presented for computing upper bounds on Lyapunov exponents and their sums, and on Lyapunov dimension, among all trajectories of a dynamical system governed by ordinary differential equations. The first approach expresses a sum of Lyapunov exponents as a time average in an augmented dynamical system and then applies methods for bounding time averages. This generalizes the work of Oeri & Goluskin (Nonlinearity 36:5378–5400, 2023), who bounded the single leading Lyapunov exponent. The second approach considers a different augmented dynamical system, where bounds on sums of Lyapunov exponents are implied by stability of certain sets, and such stability is verified using Lyapunov function methods. Both of our approaches also can be adapted to directly compute bounds on Lyapunov dimension, which in turn implies bounds on the fractal dimension of a global attractor. For systems of ordinary differential equations with polynomial right-hand sides, all of our bounding formulations lead to polynomial optimization problems with sum-of-squares constraints. These sum-of-squares problems can be solved computationally for any particular system to yield numerical bounds, provided the number of variables and degree of polynomials is not prohibitive. Most of our upper bounds are proven to be sharp under relatively weak assumptions. In the case of the polynomial optimization problems, sharpness means that upper bounds converge to the exact values as polynomial degrees are raised. Computational examples demonstrate upper bounds that are sharp to several digits, including for a six-dimensional dynamical system where sums of Lyapunov exponents are maximized on periodic orbits.
@@ 1 Introduction
The detection and characterization of chaos is a major computational challenge for many dynamical systems. Chaotic properties are often quantified via the spectrum of Lyapunov exponents (LEs), which can be computed over individual trajectories to obtain local information about that trajectory and, perhaps, global implications such as existence of hyperchaos or estimates of attractor dimension. In this paper we consider the problem of finding upper bounds on sums of LEs among all possible trajectories, without knowledge of any particular trajectories. Upper bounds on sums of LEs can be used to construct upper bounds on the fractal dimension of the global attractor. Our upper bounds are complementary to typical trajectory-based computations, which give corresponding lower bounds on sums of LEs and on Lyapunov dimension of global attractors.
The LE spectrum quantifies the asymptotic rates at which linearized perturbations in different directions grow or decay along trajectories of a dynamical system. For dynamics in $n$ dimensions there are $n$ exponents, defined in decreasing order. Partial sums of the first $k$ exponents describe the exponential rates of growth or contraction of $k$-dimensional volumes in state space, so that the first exponent governs the stretching of lines, the sum of the first two governs the growth of areas, and so on. In dissipative systems, these sums become negative beyond a certain value of $k$, reflecting the contraction of higher-dimensional volumes. One way of defining a value at which this contraction begins leads to the Lyapunov dimension, which is non-integer in general, and which provides a natural connection between dynamical stability and the fractal geometry of invariant sets such as strange attractors.
@@
with $F\in C^{1}(\Omega,\mathbb{R}^{n})$, where $C^{k}(A,B)$ denotes the space of functions mapping $A$ to $B$ that are $k$ times continuously differentiable. The state space $\Omega\subset\mathbb{R}^{n}$ may be all of $\mathbb{R}^{n}$, a subdomain of $\mathbb{R}^{n}$ or a lower-dimensional manifold embedded in $\mathbb{R}^{n}$.
In the present work, we apply existing methods for proving global properties of ODE dynamical systems to the problem of bounding LEs and their sums. These existing methods rely on finding auxiliary functions $V\in C^{1}(\Omega,\mathbb{R})$ subject to certain pointwise inequalities that, in turn, imply global statements about the dynamics. The best known examples of auxiliary functions are Lyapunov functions, whose inequality conditions imply statements about stability. (Our use of Lyapunov functions to study Lyapunov exponents is not typical, despite both objects bearing the same person’s name.) Among the lesser known types of auxiliary functions are those whose inequality conditions imply bounds on infinite-time averages. For each quantity studied in the present work, we formulate two different upper bounds: a “shifted spectrum” approach based on Lyapunov function conditions for stability, and a “sphere projection” approach based on auxiliary function conditions for bounding time averages.
The shifted spectrum approach that we present in section 2.3 uses conditions on $V$ that are similar to standard Lyapunov function conditions. In stability theory, the pointwise inequality conditions required of a Lyapunov function may be, for instance,
@@
where $DV$ denotes the gradient of $V$ with respect to the state space variable $X$. Here $F(X)\cdot DV(X)$ is the Lie derivative of $V$ with respect to the flow since the usual chain rule gives $\frac{\mathrm{d}}{\mathrm{d}t}V(X(t))=V(X(t))\cdot DV(X(t))$. The existence of $V$ satisfying 2 would imply that the set of $X\in\Omega$ where $V(X)=0$ is Lyapunov stable, meaning that trajectories remain arbitrarily close to this set if they start sufficiently close [17]. Until about 25 years ago, there was no broadly applicable method for finding $V$ subject to pointwise inequalities like those in 2. Now, however, it is often possible to construct such $V$ computationally using optimization methods based on sum-of-squares (SOS) polynomials [39, 36]. In particular, if the function $F$ defining the dynamics is polynomial in the components of $X$, and if $V$ is sought from a finite-dimensional space of polynomials, then $F\cdot DV$ is also a polynomial. Deciding whether the polynomials $V$ and $F\cdot DV$ are nonnegative on $\Omega$ is prohibitively hard in general [32], but this nonnegativity can be ensured by stronger and more tractable SOS conditions. For instance, simply requiring $V$ to be SOS, meaning that it can be represented as a sum of squares of other polynomials, would imply $V\geq 0$ on all of $\mathbb{R}^{n}$. Other SOS conditions can imply nonnegativity on $\Omega$ but not on all of $\mathbb{R}^{n}$, as explained in section 2.4.
The sphere projection approach that we present in section 2.2 uses conditions on an auxiliary function $V$ for bounding time averages. For any $\Phi\in C^{0}(\Omega,\mathbb{R})$ evaluated along a trajectory $X(t)$ with initial condition $X_{0}$, define the infinite-time average $\overline{\Phi}$ as
@@
On each $X(t)$ that remains bounded forward in time, boundedness of $V(X(t))$ implies $\overline{\frac{\mathrm{d}}{\mathrm{d}t}V}=0$, and thus $\overline{F\cdot DV}=0$. Thus, on bounded trajectories,
@@
The right-hand upper bound is useful because it does not involve individual trajectories. Since 4 holds for all bounded trajectories and all $V\in C^{1}$, one can maximize the left-hand side over $X_{0}$ and minimize the right-hand side over $V$ to find
@@
Furthermore, the inequality in 5 is an equality when $\Omega$ is a compact set in which trajectories remain [46, 3], and this equality underlies sharpness of our sphere projection approach. When $\Phi$, $V$ and $F$ are polynomial, so is the expression $\Phi+F\cdot DV$ on the right-hand side of 5. In this case the min–max problem can be relaxed into a minimization over polynomial $V$ subject to an SOS constraint, as explained in section 2.4. This constitutes an SOS optimization problem, where the optimization objective is to minimize the resulting upper bound on $\overline{\Phi}$. These ideas were applied by Oeri and Goluskin [34] to bound the single leading LE. Here we extend this approach to sums of LEs and to Lyapunov dimension.
Our two approaches are the first broadly applicable computational methods that can give convergent upper bounds on global maxima of LE sums or Lyapunov dimension. Unlike methods that rely on pointwise optimization over state space [7], our methods remain sharp when the maximizing orbits are periodic, rather than stationary. Furthermore, our methods give a systematic way to search for the maximizing trajectories themselves.
The rest of the paper is organized as follows. Section 2 defines sums of LEs and Lyapunov dimension and then describes our two general approaches to bounding these quantities. After arriving at more general optimization problems, we strengthen their constraints to obtain SOS optimization problems, then we describe how to exploit symmetries. Section 3 presents computational results from applications of our methods to two examples: the chaotic Duffing oscillator, which can be formulated as an autonomous system in 4 variables, and a 3-degree-of-freedom Hamiltonian system adapted from [35], which is an autonomous system in 6 variables. For all LE sums in both examples, we confirm that our computed bounds are sharp to several digits by finding periodic orbits on which the bounds are approximately saturated. Concluding remarks are given in section 4. For completeness, appendix A summarizes necessary material about exterior products, and appendix B proves a non-standard variant of an existence theorem for Lyapunov functions.
@@ 2 Optimization problems and SOS relaxations
This section defines the objects we seek to bound and formulates our methods for doing so. section 2.1 give expressions for LE sums and Lyapunov dimension that are useful for what follows. In section 2.2 we describe approaches for bounding these quantities based on projecting tangent space vectors onto the unit sphere. In section 2.3 we describe approaches that instead shift the spectrum of the Jacobian so that tangent vectors remain bounded. Section 2.4 describes how each formulation is relaxed to into SOS optimization problems that can be implemented numerically. Finally, section 2.5 describes how symmetries may be exploited in the optimization problems.
@@ 2.1 Tangent vectors, Lyapunov exponents and Lyapunov dimension
In this section we consider autonomous ODE systems on a domain $\mathcal{B}\subset\mathbb{R}^{n}$, just as in 1 but denoted here using lowercase $x$ and $f$,
@@
with $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$. Uppercase $X$ and $F$ and domain $\Omega$ will be reserved for associated ODEs on augmented state spaces that combine the dynamics of $x$ with dynamics on its tangent space. Trajectories are assumed to remain in $\mathcal{B}$ for all $t\geq 0$. In cases where $\mathcal{B}$ is unbounded, we further assume that each trajectory remains bounded for $t\geq 0$, meaning there is no forward-time blowup. The differentiability of $f$ and the assumption of no blowup guarantee well-posedness for all positive times, meaning there exists a differentiable flow map $\phi:\mathbb{R}_{+}\times\mathcal{B}\to\mathcal{B}$ satisfying $x(t)=\phi^{t}(x_{0})$ for every initial condition $x_{0}\in\mathcal{B}$ and all $t\geq 0$.
The LEs associated with a trajectory $x(t)$ quantify the convergence or divergence of infinitesimally nearby trajectories. Linearization of 6 around each point $x(t)$ defines the dynamics of the tangent vector $y(t)$, which evolves in the tangent space $\mathbb{R}^{n}$ according to
@@
We restrict $y_{0}$ to the unit sphere,
@@
because the tangent vector dynamics are linear in $y$ and thus are unaffected by normalization of the initial condition. Considering the separation between trajectories starting at $x_{0}$ and at $x_{0}+\varepsilon y_{0}$, Taylor expanding the flow map and using 8 gives
$|\phi^{t}(x_{0})-\phi^{t}(x_{0}+\varepsilon y_{0})|=\varepsilon|y(t)|+O(\varepsilon^{2}|y_{0}|^{2}).$ (10)
@@
$|\phi^{t}(x_{0})-\phi^{t}(x_{0}+\varepsilon y_{0})|\approx\varepsilon|y_{0}|e^{\mu(x_{0},y_{0})t}.$ (11)
If $\mu<0$ this approximation can be valid for all time. If $\mu>0$ this approximation can be valid only until the time when separation between trajectories grows too large for 10 to apply, but this time approaches infinity as $\varepsilon\to 0$. Motivated by the expectation that $|y(t)|\approx e^{\mu t}$, the corresponding LE can be defined precisely as
@@
For each trajectory $x(t)$, which is determined by its initial condition $x_{0}$, the LE $\mu$ defined by 12 generally depends on the initial tangent vector direction $y_{0}$. The Oseledets ergodic theorem guarantees that different $y_{0}$ give at most $n$ different LEs [40], the largest of which is called the leading LE. We define this by 2.1 and then give equivalent expressions in sections 2.2 and 2.3 that underlie the methods we introduce there. The literature on LEs uses various definitions that are equivalent in many settings [40].
@@
The leading LE is the average growth rate of lengths in the tangent space, maximized over the initial direction. A natural generalization is to consider growth rates of $k$-dimensional volumes in the tangent space for any $k\leq n$. From this one can define leading sums of LEs and, in turn, the full spectrum of LEs [45]. It is also possible to define the spectrum of LEs directly using singular values of $D\phi^{t}$ [40], but their sums are easier to define, and the sums are what we study in the present work. Let $|y_{1}\wedge\cdots\wedge y_{k}|$ denote the volume of the parallelepiped whose edges coincide with vectors $y_{1},\ldots,y_{k}\in\mathbb{R}^{n}$. If each $y_{i}$ is a tangent vector that evolves in time according to 7, the volume of the corresponding parallelepiped generally grows or shrinks exponentially in time, so we can define an average exponential rate analogously to 12. Maximizing this rate over all initial directions of the tangent vectors gives [45]
$M_{k}(x_{0})=\sup_{\begin{subarray}{c}y_{1}(0),\dots,y_{k}(0)\in\mathbb{R}^{n}\\ |y_{1}(0)\wedge\dots\wedge y_{k}(0)|=1\end{subarray}}\limsup_{t\to\infty}\frac{1}{t}\log{|y_{1}(t)\wedge\dots\wedge y_{k}(t)|},$ (14)
where each $y_{i}(t)$ evolves as in 7. The value $M_{k}(x_{0})$ is the maximum time-averaged growth rate of $k$-dimensional volumes in the tangent space for the trajectory $x(t)$ starting at $x_{0}$. This maximum growth rate will be attained by almost every choice of tangent vectors, under the assumptions of certain multiplicative ergodic theorems [40].
The set of all $y_{1}\wedge\cdots\wedge y_{k}$ form a vector space called the exterior product space of $k$-vectors on $\mathbb{R}^{n}$. The space of $k$-vectors on $\mathbb{R}^{n}$ has dimension $n_{k}=\tfrac{n!}{k!(n-k)!}$. We can therefore identify each $y_{1}\wedge\cdots\wedge y_{k}$ with a vector denoted by $y^{k}\in\mathbb{R}^{n_{k}}$, whose Euclidean norm $|y^{k}|$ gives the volume of the parallelepiped with edges $y_{1},\ldots,y_{k}$. For time-dependent $k$-vectors where each $y_{i}(t)$ is a tangent vector satisfying the ODE 7, the $k$-vectors satisfy a corresponding ODE (see A.4),
$\frac{\mathrm{d}}{\mathrm{d}t}y^{k}(t)=Df^{[k]}(x(t))\,y^{k}(t).$ (15)
Here $Df^{[k]}$ denotes the $k^{th}$ additive compound for the matrix $Df$, and later $Df^{(k)}$ denotes the $k^{th}$ multiplicative compound. These compounds are linear transformations acting on the space of $k$-vectors, induced by the linear transformation $Df$ on vectors, and they can be computed explicitly as $n_{k}\times n_{k}$ matrices. Exterior products and compound matrices are reviewed in appendix A. The expression 14 for $M_{k}$ can be equivalently stated using $y^{k}(t)$, resulting in the following 2.2 that we choose for $M_{k}$.
@@
For $x(t)$ evolving by the ODE 6 from initial condition $x_{0}$, the Lyapunov exponents are defined such that the sum of the $k$ leading Lyapunov exponents is
$M_{k}(x_{0})=\sup_{y^{k}_{0}\in\mathbb{S}^{{n_{k}}-1}}\limsup_{t\to\infty}\frac{1}{t}\log{|y^{k}(t)|},$ (16)
where $y^{k}(t)$ evolves by 15 from initial condition $y^{k}_{0}$.
@@
so indeed $M_{k}=\sum_{i=1}^{k}\mu_{i}$. There are equivalent ways to define $\mu_{k}$ as the fundamental object and then define the $M_{k}$ as their partial sums, but for our purposes it is simpler to define $M_{k}$ directly by 16.
For systems with attractors, a primary use of LE sums is to define and study the dimension of an attractor. 2.3 gives the Kaplan–Yorke formula [16] for the global Lyapunov dimension $d_{L}$ of an ODE system; see [18] for equivalent definitions. Hausdorff dimension is perhaps a more natural way to define the dimension of a fractal set in general. For the global attractor of a dynamical system, Hausdorff dimension can be very difficult to estimate directly, but it is bounded above by $d_{L}$ [8]. Thus, the upper bounds we produce for $d_{L}$ are upper bounds on Hausdorff dimension also.
@@
For $x(t)$ evolving in $\mathcal{B}\subset\mathbb{R}^{n}$ by the ODE 6, let $j$ be the smallest integer such that $\sup_{x_{0}\in\mathcal{B}}M_{j+1}(x_{0})<0$. The global Lyapunov dimension over $\mathcal{B}$ is
@@
For particular ODEs, the Kaplan–Yorke formula is usually evaluated numerically along a chaotic trajectory, rather than being maximized over all trajectories. The result indicates a fractal dimension of the chaotic attractor, rather than of the global attractor. Global Lyapunov dimension pertains to the latter, and the so-called Eden conjecture [9, 22] predicts that the maximum in 18 will be attained on predicts that the maximum in 18 is attained on an equilibrium or periodic orbit, not a chaotic orbit. 2.4 states the fact that $j\leq d_{L}<j+1$, which we prove for completeness, along with a restatement of the definition of $d_{L}$ that we will use in sections 2.2 and 2.3.

### v2

@@ Abstract
Two approaches are presented for computing upper bounds on Lyapunov exponents and their sums, and on the Lyapunov dimension, among all trajectories of a dynamical system governed by ordinary differential equations. The first approach expresses a sum of Lyapunov exponents as a time average in an augmented dynamical system and then applies methods for bounding time averages. This generalizes the method of Oeri & Goluskin (Nonlinearity 36:5378–5400, 2023) for bounding the single largest Lyapunov exponent. The second approach considers a different augmented dynamical system, where bounds on sums of Lyapunov exponents are implied by stability of certain sets, and such stability is verified using Lyapunov function methods. Both of our approaches also can be adapted to directly compute bounds on Lyapunov dimension, which in turn imply bounds on the fractal dimension of a global attractor. For systems of ordinary differential equations with polynomial right-hand sides, all of our bounding formulations lead to polynomial optimization problems with sum-of-squares constraints. These sum-of-squares problems can be solved computationally for a chosen system to yield numerical bounds, provided the number of variables and degree of polynomials are not prohibitive. Most of our bounding formulations are proved to be sharp under mild assumptions. In the case of the polynomial optimization problems, sharpness means that upper bounds converge to the quantities being bounded as polynomial degrees are raised. Computational examples demonstrate upper bounds that are sharp to several digits, including for a six-dimensional dynamical system where sums of Lyapunov exponents are maximized on periodic orbits.
@@ 1 Introduction
The detection and characterization of chaos is a major computational challenge for many dynamical systems. Chaotic properties are often quantified via the spectrum of Lyapunov exponents (LEs), which can be computed over individual trajectories. When combined with other knowledge about stability or local attractors, LEs can imply, for example, existence of hyperchaos or estimates of attractor dimension. In this paper we formulate upper bounds on sums of LEs, and on the related Lyapunov dimension of a global attractor. These bounds apply to all possible trajectories, yet they do not require knowing any trajectories. Such upper bounds are complementary to typical trajectory-based computations, which give corresponding lower bounds on sums of LEs and Lyapunov dimension.
The LE spectrum quantifies the asymptotic rates at which linearized perturbations in different directions grow or decay along trajectories of a dynamical system. For dynamics in $n$ dimensions there are $n$ exponents, indexed in decreasing order. The sum of the first $k$ exponents for a trajectory gives the exponential rate of expansion or contraction of $k$-dimensional volumes in state space, infinitesimally close to that trajectory. The first exponent governs the stretching of lines, the sum of the first two governs the growth of areas, and so on. In conservative systems, $n$-dimensional volumes in state space are conserved, so the sum of all $n$ LEs is zero. In dissipative systems, the LE sums are negative beyond a certain value of $k$, reflecting the contraction of higher-dimensional volumes. One way of defining a value $k$ at which this contraction begins leads to the Lyapunov dimension, which is non-integer in general, and which provides a natural connection between dynamical stability and the fractal geometry of invariant sets such as strange attractors.
@@
where each trajectory $X(t)$ remains in the state space $\Omega\subset\mathbb{R}^{n}$ for $t\geq 0$, where $\Omega$ may be all of $\mathbb{R}^{n}$, a subdomain of $\mathbb{R}^{n}$ or a lower-dimensional manifold embedded in $\mathbb{R}^{n}$. Assume that $F\in C^{1}(\Omega,\mathbb{R}^{n})$, where $C^{k}(A,B)$ denotes the space of functions mapping $A$ to $B$ that are $k$ times continuously differentiable, and when $B=\mathbb{R}$ we write $C^{k}(A)$.
The present work builds on existing methods for computing global properties of ODE dynamical systems. These methods rely on finding auxiliary functions $V\in C^{1}(\Omega)$ subject to certain pointwise inequalities on the state space that, in turn, imply statements about the dynamics. The best known examples of auxiliary functions are Lyapunov functions, whose inequality conditions imply statements about stability. (Our use of Lyapunov functions to study Lyapunov exponents is not typical, despite both objects bearing the same person’s name.) Among the lesser known types of auxiliary functions are those whose inequality conditions imply bounds on infinite-time averages. For each quantity studied in the present work, we formulate two different upper bounds: a “shifted spectrum” approach based on Lyapunov function conditions for stability, and a “sphere projection” approach based on auxiliary function conditions for bounding time averages.
The shifted spectrum approach that we present in section 2.3 uses conditions on $V$ that are similar to standard Lyapunov function conditions in stability theory. Such pointwise inequality conditions may be, for instance,
@@
where $DV$ denotes the gradient of $V$ with respect to $X$. Here $F(X)\cdot DV(X)$ is the Lie derivative of $V$ with respect to the flow since the usual chain rule gives $\frac{\mathrm{d}}{\mathrm{d}t}V(X(t))=V(X(t))\cdot DV(X(t))$. The existence of $V$ satisfying 2 would imply that the subset of $\Omega$ on which $V(X)=0$ is Lyapunov stable, meaning that trajectories remain arbitrarily close to this set if they start sufficiently close [18]. Until about 25 years ago, there was no broadly applicable method for finding $V$ subject to pointwise inequalities like those in 2. Now, however, it is often possible to construct such $V$ computationally using optimization methods based on sum-of-squares (SOS) polynomials [41, 37]. In particular, if the ODE right-hand side $F$ is polynomial in the components of $X$, and if $V$ is sought from a finite-dimensional space of polynomials, then $F\cdot DV$ also belongs to a finite-dimensional polynomial space. Deciding whether the polynomials $V$ and $F\cdot DV$ are nonnegative on $\Omega$ is prohibitively hard in general [33], but this nonnegativity can be ensured by stronger and more tractable SOS conditions. For instance, simply requiring $V$ to be SOS, meaning that it can be represented as a sum of squares of other polynomials, would imply $V\geq 0$ on all of $\mathbb{R}^{n}$. Other SOS conditions can imply nonnegativity on $\Omega$ rather than on $\mathbb{R}^{n}$, as explained in section 2.4.
The sphere projection approach that we present in section 2.2 uses conditions on an auxiliary function $V$ for bounding time averages. For any $\Phi\in C^{0}(\Omega)$ evaluated along a trajectory $X(t)$ with initial condition $X_{0}$, define the infinite-time average $\overline{\Phi}$ as
@@
On each $X(t)$ that remains bounded forward in time, boundedness of $V(X(t))$ implies $\overline{\frac{\mathrm{d}}{\mathrm{d}t}V}=0$, and thus $\overline{F\cdot DV}=0$. Therefore, on each bounded trajectory in $\Omega$,
@@
This upper bound is useful because it does not involve individual trajectories. Assuming each trajectory in $\Omega$ is bounded for $t\geq 0$, one can maximize the left-hand side over $X_{0}$ and minimize the right-hand side over $V$ to find
@@
Furthermore, the inequality in 5 is an equality when the forward-invariant set $\Omega$ is compact [50, 2], and this equality underlies sharpness of our sphere projection approach. When $\Phi$, $V$ and $F$ are polynomial, so is the expression $\Phi+F\cdot DV$ on the right-hand side of 5. In this case, the min–max problem can be relaxed into a minimization subject to SOS constraints, as explained in section 2.4. The relaxation is an SOS optimization problem whose objective is to minimize the resulting upper bound on $\overline{\Phi}$. These ideas were used by Oeri and Goluskin [34] to bound the single largest LE. Here we extend this approach to sums of LEs and to Lyapunov dimension.
Our two approaches are the first broadly applicable computational methods that can give convergent upper bounds on global maxima of LE sums or Lyapunov dimension. Unlike methods that rely on pointwise optimization over state space [8], our methods remain sharp when the maximizing orbits are periodic, rather than stationary. Furthermore, our methods give a systematic way to search for the maximizing trajectories themselves, as demonstrated in section 3.1.
The rest of the paper is organized as follows. Section 2 defines sums of LEs and Lyapunov dimension and then describes our two general approaches to bounding these quantities. After arriving at more general optimization problems, we strengthen their constraints to obtain SOS optimization problems, and then we describe how to exploit symmetries. Section 3 presents computational results from applications of our methods to two examples: the chaotic Duffing oscillator, which can be formulated as an autonomous system in 4 variables, and a 3-degree-of-freedom Hamiltonian system adapted from [36], which is an autonomous system in 6 variables. For all LE sums in both examples, we confirm that our computed bounds are sharp to several digits by finding periodic orbits on which the bounds are approximately saturated. Concluding remarks are given in section 4. For completeness, appendix A summarizes necessary material about exterior products, and appendix B proves a non-standard variant of an existence theorem for Lyapunov functions.
@@ 2 Optimization problems and SOS relaxations
This section defines the quantities we seek to bound and formulates our methods for doing so. Section 2.1 gives expressions for LE sums and Lyapunov dimension that are best suited to what follows. In section 2.2 we describe approaches for bounding these quantities based on projecting tangent space vectors onto the unit sphere. In section 2.3 we describe approaches that instead shift the spectrum of the Jacobian so that tangent vectors remain bounded. Section 2.4 describes how each formulation is relaxed into SOS optimization problems that can be implemented numerically. Finally, section 2.5 describes how symmetries may be exploited in the optimization problems.
@@ 2.1 Tangent vectors, Lyapunov exponents and Lyapunov dimension
For the remainder of this paper, the ODE system whose LEs are to be studied is denoted by
@@
which is the same as 1 but with lowercase $x$ and $f$. The domain of $x(t)$ is denoted by $\mathcal{B}\subset\mathbb{R}^{n}$, so $f\in C^{1}(\mathcal{B},\mathbb{R}^{n})$. The notation in 1 with uppercase $X$ and $F$, and domain $\Omega$, will be reserved for ODEs associated with 6 on augmented state spaces that combine the dynamics of $x$ with dynamics on its tangent space. Trajectories are assumed to remain in $\mathcal{B}$ for all $t\geq 0$, meaning $\mathcal{B}$ is forward invariant. In cases where $\mathcal{B}$ is unbounded, we further assume that each trajectory remains bounded for $t\geq 0$, meaning there is no forward-time blowup. The differentiability of $f$ and the assumption of no blowup guarantee well-posedness for all positive times, meaning there exists a differentiable flow map $\phi:\mathbb{R}_{+}\times\mathcal{B}\to\mathcal{B}$ satisfying $x(t)=\phi^{t}(x_{0})$ for every initial condition $x_{0}\in\mathcal{B}$ and all $t\geq 0$.
The LEs associated with a trajectory $x(t)$ quantify the convergence or divergence of infinitesimally nearby trajectories. Linearization of 6 around $x(t)$ defines the dynamics of the tangent vector $y(t)$, which evolves in the tangent space $\mathbb{R}^{n}$ according to
@@
We can restrict $y_{0}$ to the unit sphere,
@@
because the tangent vector dynamics are linear in $y$, so scaling $y_{0}$ simply scales $y(t)$ proportionally. Considering the separation between trajectories starting at $x_{0}$ and at $x_{0}+\varepsilon y_{0}$, Taylor expanding the flow map and using 8 and $|y_{0}|=1$ gives
$|\phi^{t}(x_{0})-\phi^{t}(x_{0}+\varepsilon y_{0})|=\varepsilon|y(t)|+O(\varepsilon^{2}).$ (10)
@@
$|\phi^{t}(x_{0})-\phi^{t}(x_{0}+\varepsilon y_{0})|\approx\varepsilon e^{\mu(x_{0},y_{0})t}.$ (11)
If $\mu<0$ this approximation can be valid for all time. If $\mu>0$ this approximation can be valid only until the separation between trajectories grows too large for 10 to apply, but this time horizon approaches infinity as $\varepsilon\to 0$. Motivated by the expectation that $|y(t)|\approx e^{\mu t}$, the corresponding LE can be defined precisely as
@@
For each trajectory $x(t)$, which is determined by its initial condition $x_{0}$, the LE $\mu$ defined by 12 generally depends on the initial tangent vector direction $y_{0}$. The Oseledets ergodic theorem guarantees that different $y_{0}$ give at most $n$ different LEs [42], the largest of which is called the leading LE. 2.1 below gives an expression for the leading LE. The literature on LEs uses various definitions that are equivalent in many settings [42].
@@
The leading LE is the average growth rate of lengths in the tangent space, maximized over the initial direction $y_{0}$. A natural generalization is to consider growth rates of $k$-dimensional volumes in the tangent space for any $k\leq n$. From this one can define leading sums of LEs and, in turn, the full spectrum of LEs [49]. Let $|y_{1}\wedge\cdots\wedge y_{k}|$ denote the volume of the parallelepiped whose edges coincide with vectors $y_{1},\ldots,y_{k}\in\mathbb{R}^{n}$. If each $y_{i}$ is a time-evolving tangent vector, the volume of the parallelepiped generally grows or shrinks exponentially in time, so we can define an average exponential rate analogously to 12. Maximizing this rate over all initial directions $y_{i}(0)$ of the tangent vectors gives [49]
$M_{k}(x_{0})=\sup_{|y_{1}\wedge\dots\wedge y_{k}|_{t=0}=1}\limsup_{t\to\infty}\frac{1}{t}\log{|y_{1}(t)\wedge\dots\wedge y_{k}(t)|},$ (14)
where each $y_{i}(t)$ evolves as in 7. The value $M_{k}(x_{0})$ is the maximum time-averaged growth rate of $k$-dimensional volumes in the tangent space for the trajectory $x(t)$ starting at $x_{0}$. This maximum growth rate will be attained by almost every choice of initial tangent vectors, provided that certain multiplicative ergodic theorems apply [42].
The set of all ‘$k$-blades’ $y_{1}\wedge\cdots\wedge y_{k}$ span a vector space called the exterior product space of $k$-vectors on $\mathbb{R}^{n}$. Properties of exterior products that we require are reviewed in appendix A. The space of $k$-vectors has dimension $C_{n}^{k}=\tfrac{n!}{k!(n-k)!}$, so we can identify each $k$-blade $y_{1}\wedge\cdots\wedge y_{k}$ with a vector denoted by $y^{k}\in\mathbb{R}^{C_{n}^{k}}$, whose Euclidean norm $|y^{k}|$ gives the volume of the parallelepiped with edges $y_{1},\ldots,y_{k}$. For time-dependent $k$-blades where each $y_{i}(t)$ is a tangent vector that satisfies the ODE 7, the corresponding ODE that $y^{k}$ satisfies is
$\frac{\mathrm{d}}{\mathrm{d}t}y^{k}(t)=Df^{[k]}(x(t))\,y^{k}(t),\qquad y^{k}(0)=y^{k}_{0}.$ (15)
Here $Df^{[k]}$ denotes the $k^{\mathrm{th}}$ additive compound matrix induced by the matrix $Df$. This compound is a linear transformation acting on the space $\mathbb{R}^{C_{n}^{k}}$ of $k$-vectors, induced by the linear transformation $Df$ on $\mathbb{R}^{n}$, and it can be computed explicitly as a $C_{n}^{k}\times C_{n}^{k}$ matrix. Additive (and multiplicative) compound matrices are reviewed in appendix A. The $y^{k}$ dynamics 15 follow from the $y$ dynamics 7 after applying the product and chain rule to $y_{1}\wedge\dots\wedge y_{k}$ and using the second part of proposition A.3. The expression 14 for $M_{k}$ can be equivalently stated using $y^{k}$, yielding the following 2.2 that we choose for $M_{k}$.
@@
For $x(t)$ evolving by the ODE 6, the Lyapunov exponents are defined such that the sum of the $k$ leading Lyapunov exponents is
$M_{k}(x_{0})=\sup_{y^{k}_{0}\in\mathbb{S}^{{C_{n}^{k}}-1}}\limsup_{t\to\infty}\frac{1}{t}\log{|y^{k}(t)|},$ (16)
where $y^{k}(t)$ evolves by 15.
@@
so indeed $M_{k}=\sum_{i=1}^{k}\mu_{i}$. There are equivalent ways to define the $\mu_{k}$ directly using singular values, and then define the $M_{k}$ as their sums [42], but for our purposes it is simpler to define $M_{k}$ directly by 16.
For systems with attractors, a primary use of LE sums is to define and study the dimension of an attractor. A notion of dimension for fractal sets in general is the Hausdorff dimension, which can be very hard to approximate for fractals that arise from dynamical systems. However, Hausdorff dimension is bounded above by Lyapunov dimension [9], which is given by the Kaplan–Yorke formula [17] in 2.3. (See [19] for equivalent definitions of Lyapunov dimension.) Our methods give upper bounds on Lyapunov dimension of global attractors, which also are upper bounds on Hausdorff dimension of global attractors.
@@
Let $x(t)$ evolve by the ODE 6, and let $\mathcal{B}$ be forward invariant. Define $j$ as the smallest integer such that $\sup_{x_{0}\in\mathcal{B}}M_{j+1}(x_{0})<0$, and define the global Lyapunov dimension over $\mathcal{B}$ as
@@
For particular ODEs, the Kaplan–Yorke formula is usually evaluated numerically along a chaotic trajectory, rather than being maximized over all trajectories. This approximates the fractal dimension of the chaotic attractor, rather than the global attractor. Global Lyapunov dimension pertains to the latter, and the so-called Eden conjecture [10, 23] predicts that the supremum in 18 is attained on an equilibrium or periodic orbit, not a chaotic orbit. 2.4 states that $j\leq d_{L}<j+1$, which we prove for completeness, and gives a rearranged formula for $d_{L}$ that we use in sections 2.2 and 2.3.